The perimeter of a rectangle is 66m. If the width were doubled and the length were increased by 22m, the perimeter would be 136m. What are the length and width of the rectangle?
Answer by checkley71(8403)

2L+2W=66 THEN 2L=66-2W AND
2(L+22)+2(2W)=136 SUBSTITUTING (66-2W) THEN
2L+44+4W=136
(66-2W)+44+4W=136
66-2W+44+4W=136
2W=136-66-44
2W=26
W=26/2
W=13 SOLUTION THEN
2L=66-2*13
2L=66-26
2L=40 SOLUTION
L=40/2
L=20
PROOF
2(20+22)+2(2*13)=136
2*42+2*26=136
84+52=136
136=136